OpenStudy (anonymous):

Evaluate for x = 2 and y = 3: x^2-y^-3/x^-1y

OpenStudy (anonymous):

Input 2 for x and 3 for y.

OpenStudy (anonymous):

\(\Large\frac{x^2-y^{-3}}{x^{-1}y}\) \(\Large\frac{(2)^2-(3)^{-3}}{(2)^{-1}(3)}\) Can you simplify @sarah841

OpenStudy (anonymous):

.3 is a rational number?

OpenStudy (anonymous):

i got 34/9 idk

OpenStudy (anonymous):

\(\Large\frac{4-0.04}{(0.5)(3)}\)

OpenStudy (anonymous):

\(\Large\frac{3.96}{1.50}\) Can you divide that? @sarah841

OpenStudy (anonymous):

2.64

OpenStudy (anonymous):

Correct. That's your answer..

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

im confused though because the answer is supposed to be 8/81 or 2/9 or 1/648

OpenStudy (anonymous):

Do you have choices?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8/81 or 2/9 or 1/648

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well..it can be the fraction 214/81..but idk..